OpenStudy (anonymous):

vertices:15 Edges:24 faces: ? 9 10 11 12

OpenStudy (wolf1728):

Does this have something to do with Euler's formula?

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

V - E + f = 2

OpenStudy (wolf1728):

15 -24 -2 = -f -11 -f f = 11

OpenStudy (anonymous):

I think it is a because u have to take 24 from 15 and u will get 9

OpenStudy (wolf1728):

I think the answer is 11

OpenStudy (wolf1728):

Remember to subtract the 2